OpenStudy (anonymous):

Please help me!!!! Which ordered pair is a solution for 4x - 3y = 6? A.(-3,2) B.(3,2) C.(3,-2) D.(-3,-2)

OpenStudy (hockeychick23):

4(-3)-3(2)= -12-6= -18 so you can eliminate A 4(3)-3(2)= 12-6=6 so its B

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

thank you guys soo much :)